6-2 A Compute the EUAB for these cash flows based on a 10% interest rate.
6000 4500 3000 1500
Answer #1

Cash inflows in the first year = 1500

Increases by 1500 every subsequent years

Life = 4 years

Interest Rate = 10%

Calculate EUAB or Annual Worth

Here, cash flow in the first year is 1500 and gradient is 1500

A = A1+G(A/G, 10%, 4)

A = 1500+ 1500(1.3812)

A = 3571.8 or rounded off to 3572

Alternatively , calculate present worth

PV = 1500(P/A, 10%, 4) + 1500(P/G, 10%, 4)

PV = 1500(3.1699)+1500(4.3781)=11322

AW= PV(A//P,10%,4)

AW =11322(0.3155)

AW = 3572

The EUAB OF THE GIVE CASH FLOWS WILL BE $3572.